# GRITS: A Spillage-Aware Guided Diffusion Policy for Robot Food Scooping Tasks

### v1

@@ Abstract
Robotic food scooping is a critical manipulation skill for food preparation and service robots. However, existing robot learning algorithms, especially learn-from-demonstration methods, still struggle to handle diverse and dynamic food states, which often results in spillage and reduced reliability. In this work, we introduce GRITS: A Spillage-Aware Guided Diffusion Policy for RobotIc Food Scoop TaskS. This framework leverages guided diffusion policy to minimize food spillage during scooping and to ensure reliable transfer of food items from the initial to the target location. Specifically, we design a spillage predictor that estimates the probability of spillage given current observation and action rollout. The predictor is trained on a simulated dataset with food spillage scenarios, constructed from four primitive shapes (spheres, cubes, cones, and cylinders) with varied physical properties such as mass, friction, and particle size. At inference time, the predictor serves as a differentiable guidance signal, steering the diffusion sampling process toward safer trajectories while preserving task success. We validate GRITS on a real-world robotic food scooping platform. GRITS is trained on six food categories and evaluated on ten unseen categories with different shapes and quantities. GRITS achieves an 82% task success rate and a 4% spillage rate, reducing spillage by over 40% compared to baselines without guidance, thereby demonstrating its effectiveness.
@@ I INTRODUCTION
In our experiments, we demonstrate the effectiveness of GRITS in successfully scooping a wide range of food items while significantly reducing spillage. We compare GRITS against several strong baselines, including rule-based methods, learning-based approaches without spillage guidance [14, 15], and variations of guided diffusion policies. Experiments are conduct on a real-world robotic platform with diverse food types varying in shape, particle size, and quantity, with six granular food categories for training and ten unseen categories for testing. The results show that GRITS generalizes well to novel scenarios, achieving an 82% success rate and the lowest spillage rate of 4%, which represents a reduction of more than 40% compared to standard diffusion policy baselines.
@@ II-A Robotic Food Manipulation
Food manipulation has attracted significant attention in the robotics community, with a growing number of studies addressing its challenges and applications, such as food preparation [1, 2, 3, 4] and assistive feeding [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Within this domain, robotic food scooping has emerged as a core task, driving research on both system design and algorithmic solutions. Recent advances in this direction focus on addressing specific challenges: developing learning frameworks to handle diverse food properties and types [5, 7, 10, 11, 14], designing specialized hardware for food acquisition [9], and creating integrated assistive feeding systems [22]. These learning-based approaches are instances of learning from demonstrations (LfD), relying on high-quality expert demonstrations [23, 24]. However, collecting demonstrations across diverse foods and conditions is costly, and existing approaches mainly optimize for task success without accounting for dynamic food states or fine-grained control, often resulting in spillage in real-world settings. In this work, we present a new LfD framework based on guided diffusion policies that accounts for unseen food states, preventing spillage while ensuring successful scooping.
@@ III-C1 Simulated Food Scooping Data Collection
We utilize Isaac Lab [20] to construct an environment with hardware settings that closely replicate the real-world setup to mitigate the sim-to-real gap. Fig. 3 depicts our simulation development environment. We construct dataset of 4,000 scooping trials, with 2,000 spillage and 2,000 non-spillage Trajectories are diversified by applying random shifts to a mean trajectory derived from real-world demonstrations, with noise ranges of ±7 cm along the X-axis and ±5 cm along the Y and Z axes. To ensure feasibility, the perturbations are constrained by both the base trajectory and bowl size. Further diversity is introduced by varying simulated object properties, including particle shape, radius, mass, friction, and quantity. The dataset spans four primitive shapes: spheres, cubes, cones, and cylinders, with 1,000 trials for each.
@@ III-C2 Spillage Predictor Training
The spillage predictor $f_{\mathtt{spillage}}$ estimates the probability of spillage $P_{\mathtt{spillage}}$} by taking a generated trajectory $A_{t}$} and a segmented object point cloud $P_{t}$}. We use segmented point clouds $P_{t}$ as the environment representation to reduce the sim-to-real gap. To further reduce the influence of irrelevant scene elements, we apply SAM2 [37] to segment pixels corresponding to spoon, bowl, and food. The food point cloud is reconstructed from the masked depth images, while the spoon and bowl point clouds are represented by their known 3D models, which remain fixed in both simulation and real-world settings. We downsample each object’s point cloud to 3,000 points, reducing computational cost while keeping a consistent input dimension across object types.
We encode object point clouds using the DP3 encoder [32], a variant of PointNet++ [38] that omits the T-Net and BatchNorm layers, which is effective for control tasks with a fixed camera and stable viewpoint. The predicted trajectory $A_{t}$} is encoded with a fully connected MLP, and its embedding is concatenated with the point cloud features. This joint representation is passed through an MLP decoder to produce a probability distribution over spillage and non-spillage classes via softmax. The model is trained with cross-entropy loss between the predicted probability $P_{\mathtt{spillage}}$} and the ground-truth label $y\in{0,1}$.
@@ III-C3 Guided Diffusion Policy with Spillage Predictor
We set a threshold $s=30$ to delay the activation of guidance, allowing the diffusion policy to first reconstruct a coarse trajectory before refinement and thus avoiding over-correction in the early denoising steps. For diffusion, we follow the real-world setup in Diffusion Policy [15] with DDIM [39], using $T=100$ steps for training and $T=16$ for evaluation. For guidance, we use 16 denoise iterations with a weight of $\rho=2.5$. These settings balance trajectory quality, inference efficiency, and refinement effectiveness.
Figure 5: Real-World Experimental Platform. We set up a 35 × 30 cm workspace (indicated in the Red bounding box) for the experiments.
@@ IV-A Experiment Setup
In the real-world setup, we use a 7-DoF Franka Emika Panda robot with a spoon attachment and two fixed Orbbec Femto Bolt RGB-D cameras, as shown in Fig. 5. One container is fixed, and the other is randomly placed within a 35 $\times$ 30 cm workspace. The robot controller operates at 10 Hz. To train the scooping policy, we collect expert demonstrations via a record-and-replay procedure [14]: a human kinesthetically guides the robot to record end-effector trajectories, which are then replayed to capture corresponding sensory data. The collected dataset includes six food categories: brown rice, soybeans, mung beans, chocolate balls, dried jujube, and lemon. For each category, three quantity levels and five bowl configurations are used, except for lemon, which has only one quantity level. A total of 80 demonstrations are collected. For evaluation, we prepare ten additional food categories, including sphere-like items and others with diverse shapes and properties. Each category is tested at two quantity levels with five trials each. Fig. 4 shows all the food categories and corresponding quantities.
@@ IV-B Baselines
We compare our approach with two baseline types: (1) methods without a spillage predictor, using predefined rules or naive behavior cloning, and (2) diffusion policy methods with different guidance mechanisms.
@@
  • SCONE [14]: A powerful scooping framework that integrates active perception to improve food understanding and generalization to unseen categories. It serves as a strong baseline by enhancing task success through perception-driven adaptability, though it does not explicitly tackle safety concerns such as spillage.
@@ IV-C Evaluation Metrics
We define the three evaluation metrics to access the performance of food scooping algorithms:
@@
Figure 8: Real-world rollout of GRITS. This figure illustrates GRITS scooping penne, where the original trajectory is refined into a safer one through spillage-aware guidance, reducing high-risk states. The robot then transfers the scooped items and places them into the target container. For more examples, please refer to our supplementary video.
@@ IV-D Real-World Experiments
Baselines without spillage predictor perform poorly. The rule-based method fails to adapt actions to different food items, underscoring the need for learning-based approaches. Behavior cloning and vanilla diffusion policy improve success but still suffer from high spillage rates, raising safety concerns. SCONE, which incorporates active perception, outperforms naive imitation learning by reducing spillage, but it suffers from higher scooping failure rates when handling very small food quantities, where interaction increases food state complexity. Adding spillage-aware guidance reduces spillage across all baselines. The diffusion policy with post-processing achieves a lower spillage rate 8%. However, it suffers from a high scoop failure rate 40% due to the lack of task-oriented guidance, resulting in low overall success. Among all methods, GRITS achieves the best balance, demonstrating strong robustness across food categories. We also observe a slightly higher scoop failure rate than the vanilla diffusion policy, reflecting a trade-off where spillage-aware guidance reduces risky trajectories but occasionally misses food collection.
@@
Figure 9: Ablation study of spillage predictor. We compare three variants of the spillage predictor: (1) using raw point cloud as input during real-world inference, (2) replacing the encoder with original PointNet++ [38], and (3) our GRITS method with composite point cloud (food from raw camera data, spoon and bowl from known object models) input and DP3 [32] encoder. The results show that GRITS achieves the highest success rate and lowest spillage by reducing noise while capturing critical features.
@@ IV-E Ablation Study of Spillage Predictor
We further evaluate the design of the spillage predictor. To reduce the sim-to-real gap, we construct composite point clouds by combining depth-reconstructed food with known object models of the spoon and bowl, which alleviates point dropouts and geometric distortion. To validate this design, we test the predictor on raw sensory point clouds. For feature extraction, we adopt the DP3 encoder [32]; in the ablation, we replace it with the original PointNet++ [38] for comparison. As shown in Fig. 9, raw sensory point clouds reduce robustness because noise and missing geometry obscure fine contact states. Similarly, PointNet++ has limited ability to detect object and motion patterns that lead to spillage. We often observe that these weaker predictors fail to raise alerts when trajectories approach high-risk states, so corrective refinements are triggered too late, sometimes only after spillage has already occurred. In contrast, the composite point cloud mitigates sensing artifacts, while DP3’s modified architecture captures features more effectively, yielding higher success rates and substantially lower spillage. Overall, these results confirm that both composite point cloud and DP3 encoding are critical for robust spillage-aware guidance.
V LIMITATIONS and FUTURE WORKS
Our experiments reveal that failure cases often arise from the lack of detailed information about the physical properties of food, beyond what can be captured by visual features alone. Attributes such as hardness, viscosity, and deformability strongly influence scooping outcomes and lead to difficulties across diverse food types. To address this limitation, future work could incorporate pre-interaction strategies [14] and multimodal sensing [12], including force-torque and tactile feedback, to build a more comprehensive representation of food characteristics and improve robustness. We also aim to extend this framework to more complex food manipulation tasks and explore its potential in personalized adaptive settings.

### v2

@@ Abstract
Robotic food scooping is a critical manipulation skill for food preparation and service robots. However, existing robot learning algorithms, especially learn-from-demonstration methods, still struggle to handle diverse and dynamic food states, which often results in spillage and reduced reliability. In this work, we introduce GRITS: A Spillage-Aware Guided Diffusion Policy for RobotIc Food Scoop TaskS. This framework leverages guided diffusion policy to minimize food spillage during scooping and to ensure reliable transfer of food items from the initial to the target location. Specifically, we design a spillage predictor that estimates the probability of spillage given current observation and action rollout. The predictor is trained on a simulated dataset with food spillage scenarios, constructed from four primitive shapes (spheres, cubes, cones, and cylinders) with varied physical properties such as mass, friction, and particle size. At inference time, the predictor serves as a differentiable guidance signal, steering the diffusion sampling process toward safer trajectories while preserving task success. We validate GRITS on a real-world robotic food scooping platform. GRITS is trained on six food categories and evaluated on ten unseen categories with different shapes and quantities. GRITS achieves an 82% task success rate and a 4% spillage rate, reducing spillage by over 40% compared to baselines without guidance, thereby demonstrating its effectiveness. More details are available on our project website: https://hcis-lab.github.io/GRITS/.
@@ I INTRODUCTION
In our experiments, we demonstrate the effectiveness of GRITS in successfully scooping a wide range of food items while significantly reducing spillage. We compare GRITS against several strong baselines, including rule-based methods, learning-based approaches without spillage guidance [32, 6], and variations of guided diffusion policies. Experiments are conducted on a real-world robotic platform with diverse food types varying in shape, particle size, and quantity, with six granular food categories for training and ten unseen categories for testing. The results show that GRITS generalizes well to novel scenarios, achieving an 82% success rate and the lowest spillage rate of 4%, which represents a reduction of more than 40% compared to standard diffusion policy baselines.
@@ II-A Robotic Food Manipulation
Food manipulation has attracted significant attention in the robotics community, with increasing work on applications such as food preparation [26, 35, 37, 39] and assistive feeding [4, 8, 9, 13, 14, 17, 18, 30, 31, 32]. Within this domain, robotic food scooping has emerged as a core task, driving research on both system design and algorithms. Recent advances focus on addressing specific challenges: developing learning frameworks to handle diverse food properties and types [4, 9, 17, 18, 32], designing specialized hardware for food acquisition [14], and creating integrated assistive feeding systems [21]. These learning-based approaches are instances of learning from demonstrations (LfD), relying on high-quality expert demonstrations [2, 3]. However, collecting demonstrations across diverse foods and conditions is costly, and existing approaches mainly optimize for task success without accounting for dynamic food states and fine-grained control, often leading to spillage in real-world scenarios. In this work, we present a new LfD framework based on guided diffusion policies that accounts for unseen food states, preventing spillage while ensuring successful scooping.
@@ III-C1 Simulated Food Scooping Data Collection
We utilize Isaac Lab [19] to construct an environment with hardware settings that closely replicate the real-world setup to mitigate the sim-to-real gap. Fig. 3 depicts our simulation development environment. We construct a dataset of 4,000 scooping trials, with 2,000 spillage and 2,000 non-spillage. Trajectories are diversified by applying random shifts to a mean trajectory derived from real-world demonstrations, with noise ranges of ±7 cm along the X-axis and ±5 cm along the Y and Z axes. To ensure feasibility, the perturbations are constrained by both the base trajectory and bowl size. Further diversity is introduced by varying simulated object properties, including particle shape, radius, mass, friction, and quantity. The dataset spans four primitive shapes: spheres, cubes, cones, and cylinders, with 1,000 trials for each.
@@ III-C2 Spillage Predictor Training
The spillage predictor $f_{\mathtt{spillage}}$ estimates the probability of spillage $P_{\mathtt{spillage}}$} by taking a generated trajectory $A_{t}$} and a segmented object point cloud $P_{t}$}. We use segmented point clouds $P_{t}$ as the environment representation to reduce the sim-to-real gap. To further reduce the influence of irrelevant scene elements, we apply SAM2 [25] to segment pixels corresponding to spoon, bowl, and food. The food point cloud is reconstructed from depth and the spoon and bowl are represented by their known CAD models, aligned via forward kinematics (for the spoon, rigidly attached to the end-effector) and a predefined calibrated pose (for the fixed bowl), and merged to form the composite point cloud. We downsample each object’s point cloud to 3,000 points, reducing computational cost while keeping consistent input dimension across object types.
We encode object point clouds using the DP3 encoder [38], a variant of PointNet++ [22] that omits the T-Net and BatchNorm layers, which is effective for control tasks with a fixed camera and stable viewpoint. The predicted trajectory $A_{t}$} is encoded with a fully connected MLP, and its embedding is concatenated with the point cloud features. This joint representation is passed through an MLP decoder to produce a probability distribution over spillage and non-spillage classes via softmax. The model is trained with cross-entropy loss between the predicted probability $P_{\mathtt{spillage}}$} and the ground-truth label $y\in\{0,1\}$.
@@ III-C3 Guided Diffusion Policy with Spillage Predictor
We set a threshold s=30s{=}30 to delay guidance activation, allowing the diffusion policy to form a coarse trajectory before refinement and preventing over-correction in early denoising steps. We adopt DDIM [28] following Diffusion Policy [6], with T=100T{=}100 training steps and T=16T{=}16 denoising steps for evaluation and guidance weight $\rho{=}2.5$. These settings balance trajectory quality, inference efficiency, and refinement effectiveness.
Figure 5: Real-World Experimental Platform. We set up a 35 × 30 cm workspace (indicated by the red bounding box) for the experiments.
@@ IV-A Experiment Setup
In the real-world setup, we use a 7-DoF Franka Emika Panda robot with a spoon attachment and two fixed Orbbec Femto Bolt RGB-D cameras, as shown in Fig. 5. One container is fixed, and the other is randomly placed within a 35 $\times$ 30 cm workspace. The robot controller operates at 10 Hz. To train the scooping policy, we collect expert demonstrations via record-and-replay procedure [32]: a human kinesthetically guides the robot to record end-effector trajectories, which are then replayed to capture corresponding sensory data. The collected dataset includes six food categories: brown rice, soybeans, mung beans, chocolate balls, dried jujube, and orange. For each category, three quantity levels and five bowl configurations are used, except for orange, which has only one quantity level. A total of 80 demonstrations are collected. For evaluation, we include ten additional food categories spanning sphere-like items and diverse shapes, each tested at two quantity levels with five trials per level. Fig. 4 shows all the food categories and corresponding quantities.
@@ IV-B Baselines
We compare against rule-based or naive cloning methods without a spillage predictor, and diffusion policies with alternative guidance mechanisms.
@@
  • SCONE [32]: A powerful scooping framework that integrates active perception to improve food understanding and generalization to unseen categories. It provides a strong baseline via perception-driven adaptability, but does not explicitly address safety issues such as spillage.
@@ IV-C Evaluation Metrics
We report three performance metrics evaluated over all trials, defined separately and potentially overlapping to capture complementary aspects of task performance:
@@
Figure 8: Real-world rollout of GRITS. The figure shows GRITS scooping penne, where spillage-aware guidance refines the original trajectory into a safer one by reducing high-risk states, followed by transferring the food into the target container.
@@ IV-D Real-World Experiments
Baselines without spillage predictor perform poorly. The rule-based method fails to adapt actions to different food items, underscoring the need for learning-based approaches. Behavior cloning and vanilla diffusion policy improve success but still suffer from high spillage rates, raising safety concerns. SCONE, which incorporates active perception, outperforms naive imitation learning by reducing spillage, but it suffers from higher scooping failure rates when handling very small food quantities, where interaction increases food state complexity. Adding spillage-aware guidance reduces spillage across all baselines. The diffusion policy with post-processing achieves a lower spillage rate of 8%. However, it suffers from a high scoop failure rate 40% due to the lack of task-oriented guidance, resulting in low overall success. Among all methods, GRITS achieves the best balance, demonstrating strong robustness across food categories. We also observe a slightly higher scoop failure rate than the vanilla diffusion policy, reflecting a trade-off where spillage-aware guidance reduces risky trajectories but occasionally misses food collection.
@@
Figure 9: Ablation study of spillage predictor. We compare three variants: (1) raw point cloud input at real-world inference, (2) encoder replaced with original PointNet++ [22], and (3) GRITS method with composite point cloud (food from raw camera data, spoon and bowl from known object models) input and DP3 [38] encoder. The results show that GRITS achieves the highest success rate and lowest spillage by suppressing noise while preserving critical features.
@@ IV-E Ablation Study of Spillage Predictor
We further study the spillage predictor design. To narrow the sim-to-real gap, we use composite point clouds that combine depth-reconstructed food with known spoon and bowl models, reducing point dropouts and geometric distortion. We validate this choice by comparing against raw sensory point clouds and by replacing the DP3 encoder [38] with original PointNet++ [22]. As shown in Fig. 9, raw point clouds suffer from noise and missing geometry, while PointNet++ struggles to capture spillage-relevant object and motion patterns. We often observe that these weaker predictors fail to raise alerts when trajectories approach high-risk states, so corrective refinements are triggered too late, sometimes only after spillage has already occurred. In contrast, composite point clouds with DP3 enable more effective feature extraction, leading to higher success rates and significantly lower spillage. Overall, both components are crucial for robust spillage-aware guidance.
V LIMITATIONS and FUTURE WORK
Our experiments reveal that failure cases often arise from insufficient information about food physical properties beyond visual cues. Future work could incorporate pre-interaction strategies [32] and multimodal sensing [30], including force-torque and tactile feedback, to build a more comprehensive representation of food characteristics. In addition, the current approach relies on known spoon and bowl models to enhance stabilization. Extending the framework to handle varying settings would improve generalization. Beyond scooping, the framework may extend to other food manipulation tasks such as cutting and pouring, with task-specific objectives and constraints left for future work.
@@ VI CONCLUSIONS
ACKNOWLEDGMENT
The work is sponsored in part by the National Science and Technology Council under grants 113-2634-F-002-007-, 114-2634-F-A49-004-, the Center for Intelligent Team Robotics and Human-Robot Collaboration under the “Top Research Centers in Taiwan Key Fields Program” of the Ministry of Education, Taiwan, and the Ministry of Education, and the Yushan Fellow Program Administrative Support Grant.